# Wasserstein Dependent Graph Attention Network for Collaborative Filtering with Uncertainty

### v1

@@ Abstract
Collaborative filtering (CF) is an essential technique in recommender systems that provides personalized recommendations by only leveraging user-item interactions. However, most CF methods represent users and items as fixed points in the latent space, lacking the ability to capture uncertainty. In this paper, we propose a novel approach, called the Wasserstein dependent Graph ATtention network (W-GAT), for collaborative filtering with uncertainty. We utilize graph attention network and Wasserstein distance to address the limitations of LightGCN and Kullback–Leibler divergence (KL) divergence to learn Gaussian embedding for each user and item. Additionally, our method incorporates Wasserstein-dependent mutual information further to increase the similarity between positive pairs and to tackle the challenges induced by KL divergence. Experimental results on three benchmark datasets show the superiority of W-GAT compared to several representative baselines. Extensive experimental analysis validates the effectiveness of W-GAT in capturing uncertainty by modeling the range of user preferences and categories associated with items.
@@ III-D2 Wasserstein dependent Mutual Information with Lipschitz Constraint
$\tau$ is the temperature parameter commonly used to control the hardness level in contrastive loss. $\sigma(\cdot)$ is the sigmoid function to maintain the 1-Lipschitz constraint.
@@ III-D3 Differences
Our work differs from [10] and other VAE based methods[21, 22, 23, 24] in several aspects. a) Fan et al. [10] incorrectly derives the relationship between contrastive loss and mutual information by $I(x_{a},x_{b})=\log\left[\frac{p(x_{a}|x_{b})}{p(x_{b})}\right]$, where factually it is $\log\left[\frac{p(x_{a}|x_{b})}{p(x_{a})}\right]$. It defines mutual information directly as the Wasserstein distance between x and y, i.e., $I_{W}(x,y)=-W_{2}(x,y)$ and also mistakenly states that $I_{W}(x,y)$ is proportional to $\log\left[\frac{p(x_{a}|x_{b})}{p(x_{b})}\right]$. Thus, it contains an incorrect mathematical derivation that leads to flawed reasoning logic. Our paper proposes a properly grounded formulation in collaborative filtering. b) Fan et al. [10] employs a sequential attention mechanism that leverages Wasserstein-based mutual information for sequential recommendation. Our proposed Wasserstein dependent attention mechanism framework is uniquely designed to address the challenges of uncertainty modeling and mutual information maximization in collaborative filtering on the bipartite graph. c) We aim to maximize the mutual information between positive pairs by optimizing $L_{WPC}$ while VAE models[21, 22, 23, 24] aims to learn a distribution representing all users and items by optimizing the reconstruction loss to fit the ideal distribution. W-GAT directly represents users and items as Gaussian distributions for similarity computation. Besides, our model sheds the decoder of the VAE model, i.e., there is no need to sample from distributions and use the reparameterization to train.
@@ III-F Complexity Analysis of W-GAT
TABLE II: Statistics of the three datasets
M1M DM MIND
#users 6039 16502 32643
#items 3628 11794 12108
#interactions 666765 126587 890241
#Categories 19 - 17
Sparsity 0.03043 0.00065 0.00225
TABLE III: Recommendation performance comparison of all considered baselines. The best performing values are boldfaced.
@@ IV-A1 Datasets
We experimented with three public real-world datasets: Movielens-1M  [^1] https://grouplens.org/datasets/movielens/, Digital_Music  [^2] https://nijianmo.github.io/amazon/index.html and MIND  [^3] https://msnews.github.io. M1M and DM are short for Movielens-1M and Digital_Music, respectively. M1M dataset is from the MovieLens website. DM dataset is from the public Amazon review dataset. MIND dataset is collected from anonymized behavior logs of the Microsoft News website. We adopt the standard 5-core preprocessing to filter out users or items with less than five interactions, and for MIND, we adopt 10-core. The detailed datasets statistics are summarized in Table II.
@@ IV-B Overall Performances (RQ1)
Based on the experimental outlined setup, we have conducted extensive evaluations that substantiate the effectiveness of W-GAT. Experimental results show significant improvements in recommendation metrics from multiple perspectives, validating the efficacy of our method. Illustrated in Table III, W-GAT consistently outperforms all the other baselines in terms of two metrics considered on both M1M and DM datasets. Especially on the DM dataset, our method exhibits an impressive improvement of 5.33% in $Recall@20$ and 10.68% in $NDCG@20$. On the MIND dataset, W-GAT outperforms the others concerning $Recall@20$ while remaining competitive with the best model on $NDCG@20$. For the other models based on Gaussian embeddings or KL divergence, they fail to achieve a competitive performance on different datasets. Especially for Graph2Gauss, the unstable outcomes show that the KL-based energy loss function is not applicable to recommendation tasks.
@@ IV-D1 Limitations of LightGCN (RQ3)
As mentioned in the previous section, LightGCN faces the challenge of unconditional propagation of variances, confusing the semantic interpretation of users and items. We conducted the following experiments to substantiate the suboptimal recommendation performance of LightGCN when propagating variance. We visualized the embeddings of two GNN encoders incorporating the BPR loss function on the M1M dataset. Figure 5 illustrates their representations by visualizing Gaussian distributions learned by the model on the two-dimensional representation by T-SNE [34]. Notably, only the mean values of the Gaussian distribution are used for better visualization. Observations reveal that unconstrained propagation of variances in LightGCN leads to clustering phenomena among users and items, depicted in Figure 5, validating the confusion of two semantics stated in Section III-B. It causes the model to recommend the closest item to a cluster of users, resulting in suboptimal performance in Table VI. The experimental settings in Table VI are the same as those in Table III. It specifically demonstrates the performance when using LightGCN and GAT to encode Gaussian distributions and training only with BPR loss.
@@ V Conclusion
In the future, more accurate ways to model the variances of users and items can be explored, and probabilistic representations are also worth exploring in terms of enhancing the diversity of recommendations.

### v2

@@ Abstract
Collaborative filtering (CF) is an essential technique in recommender systems that provides personalized recommendations by only leveraging user-item interactions. However, most CF methods represent users and items as fixed points in the latent space, lacking the ability to capture uncertainty. While probabilistic embedding is proposed to intergrate uncertainty, they suffer from several limitations when introduced to graph-based recommender systems. Graph convolutional network framework would confuse the semantic of uncertainty in the nodes, and similarity measured by Kullback–Leibler (KL) divergence suffers from degradation problem and demands an exponential number of samples. To address these challenges, we propose a novel approach, called the Wasserstein dependent Graph ATtention network (W-GAT), for collaborative filtering with uncertainty. We utilize graph attention network and Wasserstein distance to learn Gaussian embedding for each user and item. Additionally, our method incorporates Wasserstein-dependent mutual information further to increase the similarity between positive pairs. Experimental results on three benchmark datasets show the superiority of W-GAT compared to several representative baselines. Extensive experimental analysis validates the effectiveness of W-GAT in capturing uncertainty by modeling the range of user preferences and categories associated with items.
@@ III-D2 Wasserstein dependent Mutual Information with Lipschitz Constraint
$\tau$ is the temperature parameter commonly used to control the hardness level in contrastive loss. $\sigma(\cdot)$ is the sigmoid function to maintain the 1-Lipschitz constraint. Finally, the model will calculate the similarity between each user and all items, and recommend the top K most similar items to the user.
@@ III-D3 Differences
Our work differs from [10] and other VAE based methods[21, 22, 23, 24] in several aspects. a) Fan et al. [10] incorrectly derives the relationship between contrastive loss and mutual information by $I(x_{a},x_{b})=\log\left[\frac{p(x_{a}|x_{b})}{p(x_{b})}\right]$, where factually it is $\log\left[\frac{p(x_{a}|x_{b})}{p(x_{a})}\right]$. It defines mutual information directly as the Wasserstein distance between x and y, i.e., $I_{W}(x,y)=-W_{2}(x,y)$ and also mistakenly states that $I_{W}(x,y)$ is proportional to $\log\left[\frac{p(x_{a}|x_{b})}{p(x_{b})}\right]$. Thus, it contains an incorrect mathematical derivation that leads to flawed reasoning logic. Our paper proposes a properly grounded formulation in collaborative filtering. b) Fan et al. [10] employs a sequential attention mechanism that leverages Wasserstein-based mutual information for sequential recommendation. Our proposed Wasserstein dependent attention mechanism framework is uniquely designed to address the challenges of uncertainty modeling and mutual information maximization in collaborative filtering on the bipartite graph. c) We aim to maximize the mutual information between positive pairs by optimizing $L_{WPC}$ while VAE models[21, 22, 23, 24] aims to learn a distribution representing all users and items by optimizing the reconstruction loss to fit the ideal distribution. Compared to VAEs, the GCN-based framework can fully leverage the graph neural network to propagate the means and variances through nodes, thereby capturing information from neighborhoods, and subsequently modeling the user-item bipartite graph structure. W-GAT directly represents users and items as Gaussian distributions for similarity computation. Besides, our model sheds the decoder of the VAE model, i.e., there is no need to sample from distributions and use the reparameterization to train.
@@ III-E Training Details
We attach the pseudocode of W-GAT to facilitate future researchers in reproducing our work.
Algorithm 1 The procedure of W-GAT.
0:  The set of users $\mathcal{U}$; the set of items $\mathcal{I}$; the embedding size $D$; the temperature $\tau$; the loss weight $\omega$; the number of training epochs $Epochs$.
1:  Construct the user-item interaction graph $\mathcal{G}$ on the training set;
2:  Initialize Gaussian representations of users and items, $\textbf{e}_{u,i}\sim\mathcal{N}(\mu_{u,i},\Sigma_{u,i})$;
3:  for $T=1$ to $Epochs$ do
4:     Update the Gaussian representations with W-GAT according to Eq.(7).
5:     Obtain the predicted score $f(u,i)$ according to Eq.(13).
6:     Obtain the loss function $L$ according to Eq.(14).
7:     Update the parameters to minimize $L$.
8:  end for
8:  The overall refined user/item representations $\textbf{e}_{u}^{*}$, $\textbf{e}_{i}^{*}$.
@@ III-F Complexity Analysis of W-GAT
TABLE II: Recommendation performance comparison of all considered baselines. The best performing values are boldfaced.
@@ IV-A1 Datasets
We experimented with three public real-world datasets: Movielens-1M  [^1] https://grouplens.org/datasets/movielens/, Digital_Music  [^2] https://nijianmo.github.io/amazon/index.html and MIND  [^3] https://msnews.github.io. M1M and DM are short for Movielens-1M and Digital_Music, respectively. M1M dataset is from the MovieLens website. DM dataset is from the public Amazon review dataset. MIND dataset is collected from anonymized behavior logs of the Microsoft News website. We adopt the standard 5-core preprocessing to filter out users or items with less than five interactions, and for MIND, we adopt 10-core. The detailed datasets statistics are summarized in Table III.
TABLE III: Statistics of the three datasets
M1M DM MIND
#users 6039 16502 32643
#items 3628 11794 12108
#interactions 666765 126587 890241
#Categories 19 - 17
Sparsity 0.03043 0.00065 0.00225
@@ IV-B Overall Performances (RQ1)
Based on the experimental outlined setup, we have conducted extensive evaluations that substantiate the effectiveness of W-GAT. Experimental results show significant improvements in recommendation metrics from multiple perspectives, validating the efficacy of our method. Illustrated in Table II, W-GAT consistently outperforms all the other baselines in terms of two metrics considered on both M1M and DM datasets. Especially on the DM dataset, our method exhibits an impressive improvement of 5.33% in $Recall@20$ and 10.68% in $NDCG@20$. On the MIND dataset, W-GAT outperforms the others concerning $Recall@20$ while remaining competitive with the best model on $NDCG@20$. For the other models based on Gaussian embeddings or KL divergence, they fail to achieve a competitive performance on different datasets. Especially for Graph2Gauss, the unstable outcomes show that the KL-based energy loss function is not applicable to recommendation tasks.
@@ IV-D1 Limitations of LightGCN (RQ3)
As mentioned in the previous section, LightGCN faces the challenge of unconditional propagation of variances, confusing the semantic interpretation of users and items. We conducted the following experiments to substantiate the suboptimal recommendation performance of LightGCN when propagating variance. We visualized the embeddings of two GNN encoders incorporating the BPR loss function on the M1M dataset. Figure 4 illustrates their representations by visualizing Gaussian distributions learned by the model on the two-dimensional representation by T-SNE [34]. Notably, only the mean values of the Gaussian distribution are used for better visualization. Observations reveal that unconstrained propagation of variances in LightGCN leads to clustering phenomena among users and items, depicted in Figure 4, validating the confusion of two semantics stated in Section III-B. It causes the model to recommend the closest item to a cluster of users, resulting in suboptimal performance in Table VI. The experimental settings in Table VI are the same as those in Table II. It specifically demonstrates the performance when using LightGCN and GAT to encode Gaussian distributions and training only with BPR loss.
@@ V Conclusion
In the future, more accurate ways to model the variances of users and items can be explored, and probabilistic representations are also worth exploring in terms of enhancing the diversity of recommendations. In addition, we can also apply the idea of probabilistic representation to hypergraph recommendation systems, a field with great research prospects.